FEL Upgrade Project Weekly Brief
October 22-26, 2001

During FEL operations this week we continued our measurements of high efficiency (> 1%) wiggler operation and further
characterized the optical beam quality by making beam profile measurements (M2) as a function of power

We completed the project monthly reports (technical and cost-performance) and distributed them to the ONR and DOE contract

Preparations continued for the Nov. 8-9 semiannual review of the FEL Upgrade project.

Next week various members of the FEL team will be giving presentations at major national conferences:
George Neil will giving an overview of the IR Demo project at the APS Division of Plasma Physics conference in Los Angeles; Fred
Dylla will be giving a paper on materials processing applications of the FEL at the International Vacuum Congress in San Francisco,
and Michelle Shinn will be giving a status report on the FEL program at the Directed Energy Professional Society Symposium in

We had two visitors today from Dahlgren who are working on cryogenic systems for the “All Electric Ship”, Dr. E. Michael Golda,
Head of the Electrical Systems Department and Thomas Fikse, Lead Engineer.

WBS 3 (Beam Physics):
A description of the Revision 1.1.2 (hopefully final) Driver Design was completed and is posted at

GX and GQ magnet design analyses were reviewed; though the behaviors were largely acceptable, an inboard field roll-off of order
0.03% observed in both magnets is of some concern and is being examined more extensively.

Correction and diagnostic configurations for the linac-to-endloop match were considered; magnetic measurements were discussed.

WBS 4 (Injector):
A stainless steel reference electrode was treated in the newly re-commissioned ion implantation system at W&M.  The drawings for
the cesiator and charge collection ring in the new electrode are in sign-off.
At the end of this week’s FEL operations which saw a fair amount of high charge operation, we recesiated the electrode.

WBS 6 (RF):
Zone 3 - Front panel over filament boards was installed.  Floor plates installed over waveguide blocks.  Conformable cables were
ordered.  General cleanup of the area was done.

Zone 4 - General cleanup of the area was done.

Injector RF - The 3rd 100 kW klystron has been tested and approved for shipment.  It should arrive in about a week.  The klystron lift
plate drawing is signed, built, and tested.  The first klystron will be installed in its solenoid by Tuesday.  The LCW manifolds were
received for the circulators, RF loads, and HOM filters.  Weekly meetings have started on the upgrade of the HVPS and klystron

WBS 8 (Instrumentation):
A printer has been installed for use in printing barcode labels.  The print function is working and the program to link the data bases is
progressing nicely.

The first Beamviewer Camera/Lamp power interface box has been assembled and installed on one viewer.  It looks very sharp and
will definitely enhance the appearance of the upgraded machine as well as continue its role in the functionality.

The Laser web site upgrades continue.  A new format is being incorporated that will included photos from the Photo Gallery on the
home page so people will be able to see some of what we are doing as soon as they log on.  A page will also be added for data entry in
conjunction with the bar code program so labels can be printed easily from any computer with access to Laser.  The task list continues
to become more functional as well as minor improvements are made to make it more user friendly.

A circuit for control of the remote controlled camera platforms has been developed and is entering the testing and documentation

Development and evaluation of a vacuum control system for SNS continues with the evaluation of several controllers and the receipt
of a newly designed capacitor for use in the current Ion Pump Power Supply design, with slight modifications.

WBS 9 (Transport):
Optical Chicane Dipoles (GW)
•   We plan to inspect the first coil potting in three weeks.
Injector Dipoles (DU/DV)
•   DULY is revising small portions of the final assemblies that need modification.
Arc Dipoles (GY, GX, GQ)
•   Design continued on backchecking the detail drawings of the bend and Reverse Bend Dipoles (GX & GQ) with the
    comments and corrections.  Effort concentrated on bringing the faces and the gap in conformance with the magnetic
•   Dave Douglas is looking at the results of the magnetic model in anticipation of approving them.
•   The Procurement Department sent the package for the 180° Dipole (DY) out to vendors in the Department’s first
    totally electronic solicitation.  The drawings and specifications are in a WEB site available to all vendors for
    download.  This allows a better search for best value because of the expanded number of vendors allowed to
    participate.  We eliminated sending the cumbersome 3/4 inch stack of drawings and 3/4 inch pack of specifications
    and schedules.
Magnet Measurement
•   The drawings for the measurement stands for the two quadrupoles are in a procurement cycle.
•   In design of the dipole measurement stand we clarified the results of the magnetic model with AES to optimize
    design of the probe for the GX dipole’s UV branch.
•   We continued work on the probe that will measure the sextupole and the octupole.
3 inch quad (QX)
•   Al Guerra inspected the machining of the first core at New England Technicoil.  He resolved an interpretation of the
    drawings on the exact nature of the pole tip shape.
Trim Quad (QT)
•   The proposals from the quick re-bid cycle for this magnet are being evaluated
Sextupole (SF)
•   We have decided to go to DULY Research to get the 3D magnetic model of this magnet advanced.  Robin Wines
    has other duties that prevent her from making much headway.
Octupole (OT)
• Work on this magnet is on furlough.
Corrector Dipoles (DB, DJ)
• Milhous Co. continues work on the remainder of the batch.
Beam Line and Vacuum
•   Dave Waldman is working on the task order for design of the Arc Chambers utilizing the Engineering Services
    contract with Advanced Energy Systems.
•   Design work on girders and stands continued at a high level on the quadrupole telescope in front of the first arc and
    the region between the cryomodules.  A new stand concept using three posts out of one base for these congested
    areas is being incorporated in all layouts.
• Design of the LCW system started up again with presentation of the layout to the Monday Design Meeting.

WBS 10 (Wiggler):
We continue measurements of the wiggler and analysis of those measurements.  We hope to carry out the final fitup of the vacuum
chamber next week.

WBS 11 (Optics):
We had a service engineer from Veeco here to replace the focusing lens and zoom motor on the Wyko laser interferometer.  We also
received some additional training with the software.  We are continuing to detail components of the optical cavity, as well as continuing
design efforts on the water-cooled insertable mirror for the beam dump.  We made preparations to have an ultraviewer assembled next
week.  We are preparing for the upcoming semiannual review on Nov. 8.

The following work was done to support operations:

We provided a crystal rotator for the X-ray experiment.  We provided support for our FEL experiments, specifically, measuring the
beam quality as a function of power using diagnostics installed in the vault, installing a 3 micron high reflector as the outcoupler in order
to enable second harmonic lasing, measuring the polarization of the FEL 2nd harmonic laser emission, and measuring the pulsewidth of
the laser when operating at the fundamental (high efficiency lasing).  We also tried operating the laser at high intracavity power in an
attempt to create laser-induced damage to the coatings.  We may have done it; we saw a spike in the pressure near both mirrors, and
an increase in outcoupling.  More information on this will be in next week's report.

In preparation for upcoming experiments, we had a 90% OC with a calcium fluoride substrate brazed and metrology conducted on it.
We prepared a mirror set for broadband lasing (5 - 6.5) microns, and for 1.6 microns.  A 2 micron HR mirror will be brazed into a
holder next week.

We ran the full week for operations and performed several studies on the FEL.  We ran one additional day of nanotube production
studies for NASA/College of William and Mary.  That was followed by studies of mirror thermal distortion as a continuation of work
last summer.  The results are under analysis but it appears the mode movement which leads to an increase of M^2 is a function of
power and actually initially decreases as the power is raised. We spent one day on second harmonic lasing of the FEL as a
continuation of work from last January in hopes of providing a better explanation of the generation mechanism.  We uncovered further
mysteries:  the donut mode was found to have linear polarization instead of the expected radial.  We then spent two days on studies
aimed toward high FEL efficiency.  We got results which clearly showed turbulent behavior but we could not duplicate the very high
efficiencies seen by the Japanese.  Finally we attempted to find the damage limits of the FEL mirrors and increased the cavity power
in a very high Q setup.  The FEL continued lasing under all conditions tried and the mirrors are presently being removed to see if minor
damage occurred.